which equation describes a line that passes through the points (3,5) and (2-,3)
Answer by nerdybill(7384)

which equation describes a line that passes through the points (3,5) and (2,-3) (I'll assume you had a typo in the second point)
.
First, determine the slope:
m = (y2-y1)/(x2-x1)
m = (-3-5)/(2-3)
m = (-8)/(-1)
m = 8
.
Then using the slope above and one of the two given points (3,5) plug into the "point-slope" form:
y - y1 = m(x - x1)
y - 5 = 8(x - 3)
y - 5 = 8x - 24
y = 8x - 19 (this is what they're looking for)